kaitlyn and brittney are driving back to college after spring break. they plan to drive 120 miles on the first day. they would like to add the same amount to their driving distance each day. therefore, each day they will drive more than the previous day. they want to complete the 1200 mile trip in 5 days. find the number of miles they will drive on each of the five days.
Answer by ankor@dixie-net.com(22740)

kaitlyn and brittney are driving back to college after spring break.
they plan to drive 120 miles on the first day.
they would like to add the same amount to their driving distance each day.
therefore, each day they will drive more than the previous day.
they want to complete the 1200 mile trip in 5 days.
find the number of miles they will drive on each of the five days.
:
Let x = amt added to each days drive
120 + (x+120) + (2x+120) + (3x+120) + (4x+120) = 1200
5(120) + 10x = 1200
10x = 1200 - 600
10x = 600
x = 60 miles added to each days drive
:
Day 1: 120 mi
Day 2: 180
Day 3: 240
Day 4: 300
Day 5: 360
-----------
total:1200 mi